# Oxford OX1 3RH, UK

### v2

@@ 4.2 Model parameters
      INEL is the case where the nucleus is left in an excited state after the rescattering. This category only contains the situation where a single additional nucleon is emitted/knocked-out after rescattering. Since it does not affect the number of pions produced, it will not convert an event from a pion-less topology to a pion-production topology. The effects on nucleons are two-fold. Firstly, it can alter the number of signal events within each event topology. If the inelastic rescattering leads to two low-momentum protons below the detection threshold as opposed to a high-momentum proton, this signal event will be discarded as no protons are observed. Secondly, INEL invariably changes the kinematics of the rescattering particle. Be it the leading proton or the leading pion, based on which the TKI observables are calculated, the TKI distribution shape will be affected. Hence, while $S_{\textrm{INEL}}^{\textrm{N}}$ would affect all four data sets, $S_{\textrm{INEL}}^{\pi}$ would only affect T2K-$\pi^{+}$ and MINERvA-$\pi^{0}$.
@@
5 The first TKI-driven genie tunes
@@
5.1 Reduced tune: G24-c
To assess the tune’s quality, we plot predictions using G24-c for $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ in Figs. 9 and 9, respectively. For T2K-$0\pi$, T2K-$\pi^{+}$ and MINERvA-$0\pi$, the new $\chi^{2}$ values are comparable to those obtained with G24-0, changing less than the number of bins. For MINERvA-$\pi^{0}$, G24-c is distinctly better than G24-0, thereby demonstrating the possibility of simultaneous good descriptions of both pion-less and pion production samples with constrained parameters from cross-topology TKI tuning.
@@ 5.1 Reduced tune: G24-c
In MINERvA-$\pi^{0}$, the number of $\pi^{0}$s undergoing no FSI (“No Interaction”, as shown in the figures) is adjustable via the $S_{\lambda}^{\pi^{0}}$ parameter. A decrease in $S_{\lambda}^{\pi^{0}}$ reduces the size of the “No Interaction” events, as is shown in Fig. 10b when compared to Fig. 10a. Meanwhile, an increase in $\pi^{0}$ rescattering only manifests in pion-less measurements through ABS. There is indeed an increase of ABS for $\pi^{0}$ in T2K-$0\pi$ and MINERvA-$0\pi$, but the fraction is small to begin with, so the impact is small. The increase of $\pi^{0}$ rescattering can increase T2K-$\pi^{+}$ via CEX as discussed below. However, due to the significant suppression of CEX, its impact on T2K-$\pi^{+}$ is also minimal (for detailed breakdown, see Figs. 15-17 in Appendix B).
@@
As shown in Fig. 10a, the MINERvA-$\pi^{0}$ prediction from the nominal tune has considerable contributions from CEX controlled by $S_{\textrm{CEX}}^{\pi}$. It does not impact T2K-$0\pi$ and MINERvA-$0\pi$ as CEX only changes the pion type without removing them. Hence, events with pions in the final state would be rejected in these two data sets regardless of the pion charge. In principle, CEX will also migrate events between the signal and background definitions for T2K-$\pi^{+}$ when a $\pi^{0}$ is converted to a $\pi^{+}$ and vice verse. However, T2K-$\pi^{+}$ does not have a considerable CEX fraction to begin with, so changing CEX will have a minimal impact on its prediction. Hence, this is one of the most effective paths to decrease the MINERvA-$\pi^{0}$ cross section prediction without affecting other measurements, and indeed the $S_{\textrm{CEX}}^{\pi}$ is heavily suppressed as shown in Table 5 and illustrated in Fig. 10b compared to Fig. 10a. However, due to the intricate correlation between the FSI fates in the hA implementation discussed in Sec. 4.2, although $S_{\textrm{INEL}}^{\pi}$ and $S_{\textrm{PIPD}}^{\pi}$ are not modified explicitly, a considerable increase is observed in both INEL and PIPD for MINERvA-$\pi^{0}$.
@@
Figure 11: Similar to Fig. 10 but for $p_{\textrm{N}}$ with proton FSI fate decomposition. Note that the “Absorption” of the proton is referring to the absorption of the $\pi^{+}$ from the decay of $\Delta^{++}$, which could lead to emission of nucleons as discussed in Sec. 4.2.
@@
5.2 Alternative tune: G24-t
A similar improvement is achieved by the intermediate tune, G24-t, illustrated in Fig. 12. Instead of significantly increasing proton PIPD to elevate the $p_{\textrm{N}}$ tail, this tune notably enhances $R_{\textrm{SRC}}$—less-energetic final states are present and therefore an increased RES/DIS (Fig. 12a) ratio. Rather than suppressing $S_{\textrm{CEX}}^{\pi}$ (cf. Fig. 17d in Appendix B), this tune significantly increases it (Fig. 12b) to offset the further reduction of “No Interaction” $\pi^{0}$ events (smaller $S_{\lambda}^{\pi^{0}}$ in AllPar, cf. Table 5). Overall, this leads to a reduction of $75.31$ in $\chi^{2}$ (Table 5), similar to G24-c, as well as a good data-MC agreement across all data sets.
@@ 5.2 Alternative tune: G24-t
5.3 Discussion
For both G24-c and G24-t the fit results suggest extreme values, but the effect of the parameter change is less effective than the change in the parameter suggests. For the individual FSI fate cross section, due to the re-normalization step after scaling, the effective change for one particular fate is smaller than the scaling parameter magnitude as explained in Sec. 4.2. As for the total FSI cross sections, only that of $\pi^{0}$ is changed. The total $\pi^{0}$-C scattering cross section for pion kinetic energy is plotted in Fig. 13 to assess the overall impact. The shape remains similar and consistent with Fig. 2.23 ($\pi^{+}$-C reactions) in Ref. Andreopoulos:2015wxa while the peak magnitude increases from $380$ to $538$ mb, a $48\%$ increase, considerably less than the scaling factor, $S_{\lambda}^{\pi^{0}}=0.22$, seemingly suggests. Moreover, the default $\pi^{0}$ parameter values are calculated from charged pion data assuming isospin symmetry rather than extracted directly from experimental data. Hence, this modification does not violate any existing agreement with hadron scattering data.
@@ 5.3 Discussion
In addition, agreement of this tune with most of the non-TKI neutrino datasets available is unchanged before and after the tune, thereby demonstrating the physicality of the tune. In conclusion, this tune is a valid and effective model that can be used as a starting point for an analysis. Further refinement will be available once more data will be included in the fit, including non-TKI observables.
6 Summary and outlook
@@
This work represents the first global tuning effort on TKI data. Our partial tune of the SF-CFG and hA models, G24_20i_06_22c (G24-c), provides an effective theory to better describe both the neutrino-hydrocarbon pionless and pion production. The largest change in the model is demanded by the MINERvA-$\pi^{0}$ TKI measurement MINERvA:2020anu that was significantly overestimated in genie. The improvement is crucial for future precision GeV neutrino experiments. This tuning configuration has been integrated into the master branch of genie and is slated for inclusion in the upcoming release.
@@
To develop an effective model, we focused on the most sensitive parameters, $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$, of the SF-CFG and hA models, reducing the total number of parameters from $14$ to $6$. The corresponding best combination of observables is $\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if no $p_{\textrm{N}}$ is available), and $\delta p_{\textrm{TT}}$, from pionless and pion-production measurements in both T2K and MINERvA. In doing so, the pion production model is held fixed such that the tuned values of $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ depart from the established constraints from hadron measurements. A different pion production model might compensate part of these model tuning effects Yan:2024kkg . We also chose the hA FSI model for practical reasons. Its sequential treatment of MFP and rescattering should be considered as a numerical advantage (as was also pointed out in Ref. GENIE:2022qrc ). Other more sophisticated models could be considered in future tunes. With the existing four TKI measurements from T2K and MINERvA, we have also derived other effective tunes like G24_20i_fullpara_alt (G24-t). The degeneracy can be resolved with additional data, particularly from argon-based measurements MicroBooNE:2022emb ; MicroBooNE:2023cmw ; MicroBooNE:2023tzj ; MicroBooNE:2023wzy ; MicroBooNE:2024tmp ; MicroBooNE:2015bmn .
@@ Acknowledgment
Appendix A Combinations of observables
Combi- 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26
T2K-$0\pi$
T2K-$\pi^{+}$
MINERvA-$0\pi$
MINERvA-$\pi^{0}$

### v3

@@ IV.2 Model parameters
      INEL is the case where the nucleus is left in an excited state after the rescattering. This category only contains the situation where a single additional nucleon is emitted/knocked-out after rescattering. Since it does not affect the number of pions produced, it will not convert an event from a pionless topology to a pion-production topology. The effects on nucleons are two-fold. Firstly, it can alter the number of signal events within each event topology. If the inelastic rescattering leads to two low-momentum protons below the detection threshold as opposed to a high-momentum proton, this signal event will be discarded as no protons are observed. Secondly, INEL invariably changes the kinematics of the rescattering particle. Be it the leading proton or the leading pion, based on which the TKI observables are calculated, the TKI distribution shape will be affected. Hence, while $S_{\textrm{INEL}}^{\textrm{N}}$ would affect all four data sets, $S_{\textrm{INEL}}^{\pi}$ would only affect T2K-$\pi^{+}$ and MINERvA-$\pi^{0}$.
@@
V The first TKI-driven genie tunes
@@
V.1 Reduced tune: G24-c
To assess the tune’s quality, we plot predictions using G24-c for $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ in Figs. 9 and 9, respectively. For T2K-$0\pi$, T2K-$\pi^{+}$ and MINERvA-$0\pi$, the new $\chi^{2}$ values are comparable to those obtained with G24-0, changing less than the number of bins. For MINERvA-$\pi^{0}$, G24-c is distinctly better than G24-0, thereby demonstrating the possibility of simultaneous good descriptions of both pionless and pion production samples with constrained parameters from cross-topology TKI tuning. Note that in the region of $p_{\textrm{N}}\sim 0.3~{}\textrm{GeV}/c$, the model deficit previously reported in Ref. [20] (and also shown in Fig. 3c) persists even after the fit. The physical origin of this deficit might be attributed to the strength of the 2p2h contribution [20], but it is still a topic of significant discussion within the community.
@@ V.1 Reduced tune: G24-c
In MINERvA-$\pi^{0}$, the number of $\pi^{0}$s undergoing no FSI (“No Interaction”, as shown in the figures) is adjustable via the $S_{\lambda}^{\pi^{0}}$ parameter. A decrease in $S_{\lambda}^{\pi^{0}}$ reduces the size of the “No Interaction” events, as is shown in Fig. 10b when compared to Fig. 10a. Meanwhile, an increase in $\pi^{0}$ rescattering only manifests in pionless measurements through ABS. There is indeed an increase of ABS for $\pi^{0}$ in T2K-$0\pi$ and MINERvA-$0\pi$, but the fraction is small to begin with, so the impact is small. The increase of $\pi^{0}$ rescattering can increase T2K-$\pi^{+}$ via CEX as discussed below. However, due to the significant suppression of CEX, its impact on T2K-$\pi^{+}$ is also minimal (for detailed breakdown, see Figs. 15-17 in Appendix .2).
@@
As shown in Fig. 10a, the MINERvA-$\pi^{0}$ prediction from the nominal tune has considerable contributions from CEX controlled by $S_{\textrm{CEX}}^{\pi}$. It does not impact T2K-$0\pi$ and MINERvA-$0\pi$ as CEX only changes the pion type without removing them. Hence, events with pions in the final state would be rejected in these two data sets regardless of the pion charge. In principle, CEX will also migrate events between the signal and background definitions for T2K-$\pi^{+}$ when a $\pi^{0}$ is converted to a $\pi^{+}$ and vice verse. However, T2K-$\pi^{+}$ does not have a considerable CEX fraction to begin with, so changing CEX will have a minimal impact on its prediction. Hence, this is one of the most effective paths to decrease the MINERvA-$\pi^{0}$ cross section prediction without affecting other measurements, and indeed the $S_{\textrm{CEX}}^{\pi}$ is heavily suppressed as shown in Table 5 and illustrated in Fig. 10b compared to Fig. 10a. However, due to the intricate correlation between the FSI fates in the hA implementation discussed in Sec. IV.2, although $S_{\textrm{INEL}}^{\pi}$ and $S_{\textrm{PIPD}}^{\pi}$ are not modified explicitly, a considerable increase is observed in both INEL and PIPD for MINERvA-$\pi^{0}$.
Another impact of the larger $S_{\textrm{PIPD}}^{\textrm{N}}$ manifests in the appearance of 2p2h ($1\%$) and CCQE ($2\%$) in the MINERvA-$\pi^{0}$ cross section, as shown in Fig. 9d. These neutrino-nucleon interactions do not produce pions, but the product nucleons can produce pions via re-scattering in the nucleus, leading to a contribution to topologies with pions in the final states. However, such contribution remains small.
@@
Figure 11: Similar to Fig. 10 but for $p_{\textrm{N}}$ with proton FSI fate decomposition. Note that the “Absorption” of the proton is referring to the absorption of the $\pi^{+}$ from the decay of $\Delta^{++}$, which could lead to emission of nucleons as discussed in Sec. IV.2.
Although the interaction models, e.g. 2p2h, are not directly tuned, the respective fractions have undergone small changes in Fig. 9 and Fig. 9, compared to Fig. 3 and Fig. 3. This is a result of the increased $R_{\textrm{SRC}}$ which leads to more events with higher initial nucleon momenta such that interactions requiring higher initial state energy, such as RES and DIS, will be more frequent. This is reflected by the increase in the ratios of these interactions: For example, RES increases from $19\%$ (Fig. 3c) to $25\%$ (Fig. 9c) for MINERvA-$0\pi$. Hence, the other interactions, although not directly tuned, will experience a decrease in the fraction of total cross section: For example, 2p2h drops from $34\%$ (Fig. 3c) to $30\%$ (Fig. 9c) for MINERvA-$0\pi$.
V.2 Alternative tune: G24-t
A similar improvement is achieved by the intermediate tune, G24-t, illustrated in Fig. 12. Instead of significantly increasing proton PIPD to elevate the $p_{\textrm{N}}$ tail, this tune notably enhances $R_{\textrm{SRC}}$—more-energetic nucleons are present and therefore the RES $p_{\textrm{N}}$ peak is shifted to the right in Fig. 12a with respect to Fig. 9d. Rather than suppressing $S_{\textrm{CEX}}^{\pi}$ (cf. Fig. 17d in Appendix .2), this tune significantly increases it (Fig. 12b) to offset the further reduction of “No Interaction” $\pi^{0}$ events (smaller $S_{\lambda}^{\pi^{0}}$ in AllPar, cf. Table 5). Overall, this leads to a reduction of $75.31$ in $\chi^{2}$ (Table 5), similar to G24-c, as well as a good data-MC agreement across all data sets.
@@ V.2 Alternative tune: G24-t
V.3 Discussion
For both G24-c and G24-t the fit results suggest extreme values, but the effect of the parameter change is less effective than the change in the parameter suggests. For the individual FSI fate cross section, due to the re-normalization step after scaling, the effective change for one particular fate is smaller than the scaling parameter magnitude as explained in Sec. IV.2. As for the total FSI cross sections, only that of $\pi^{0}$ is changed. The total $\pi^{0}$-C scattering cross section for pion kinetic energy is plotted in Fig. 13 to assess the overall impact. The shape remains similar and consistent with Fig. 2.23 ($\pi^{+}$-C reactions) in Ref. [52] while the peak magnitude increases from $380$ to $538$ mb, a $48\%$ increase, considerably less than the scaling factor, $S_{\lambda}^{\pi^{0}}=0.22$, seemingly suggests. Moreover, the default $\pi^{0}$ parameter values are calculated from charged pion data assuming isospin symmetry rather than extracted directly from experimental data. Hence, this modification does not violate any existing agreement with hadron scattering data.
@@ V.3 Discussion
In addition, agreement of this tune with most of the non-TKI neutrino datasets available is unchanged before and after the tune, thereby demonstrating the physicality of the tune.
Exclusive electron scattering data is better suited to constrain FSI models. Regrettably, most electron scattering data to date is inclusive [64] and it is not well suited for the tuning of FSI. The “Electrons for Neutrinos” (e4nu) collaboration exploits data from the CLAS6 spectrometer to measure exclusive final states, such as a $1\textrm{p}0\pi$ cross-section measurement on carbon as a function of TKI variables [28]. But for electron-scattering pion production, currently, the genie predictions over-predict data due to the lack of tuning of the genie models on free nucleon data. Such big uncertainties in the pion production model mask possible nuclear model or FSI dependence. In order to be sensitive to these effects, one need to first tune electron-scattering to free nucleon data. The current tuning study could be extended in the future when data from dedicated electro-pion production measurements are available.
In conclusion, this tune is a valid and effective model that can be used as a starting point for an analysis. Further refinement will be available once more data will be included in the fit, including non-TKI observables and electron-scattering data.
VI Summary and outlook
@@
This work represents the first global tuning effort on TKI data. Our partial tune of the SF-CFG and hA models, G24_20i_06_22c (G24-c), provides an effective theory to better describe both the neutrino-hydrocarbon pionless and pion production. The largest change in the model is demanded by the MINERvA-$\pi^{0}$ TKI measurement [14] that was significantly overestimated in genie. The improvement is crucial for future precision GeV neutrino experiments. This tuning configuration has been integrated into the master branch of genie and is slated for inclusion in the upcoming release.
@@
To develop an effective model, we focused on the most sensitive parameters, $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$, of the SF-CFG and hA models, reducing the total number of parameters from $14$ to $6$. The corresponding best combination of observables is $\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if no $p_{\textrm{N}}$ is available), and $\delta p_{\textrm{TT}}$, from pionless and pion-production measurements in both T2K and MINERvA. In doing so, the pion production model is held fixed such that the tuned values of $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ depart from the established constraints from hadron measurements. A different pion production model might compensate part of these model tuning effects [65]. We also chose the hA FSI model for practical reasons. Its sequential treatment of MFP and rescattering should be considered as a numerical advantage (as was also pointed out in Ref. [11]). Other more sophisticated models could be considered in future tunes. With the existing four TKI measurements from T2K and MINERvA, we have also derived other effective tunes like G24_20i_fullpara_alt (G24-t). The degeneracy can be resolved with additional data, particularly from argon-based measurements [23, 24, 25, 26, 27, 66] and e4nu data [28].
@@ Acknowledgment
appendix
.1 Combinations of Observables
@@
Combi- 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 (Best- AllPar) 16 17 18 19 20 21 22 23 24 (Super- set) 25 26 (Best- RedPar)
T2K-$0\pi$
T2K-$\pi^{+}$
MINERvA-$0\pi$
MINERvA-$\pi^{0}$